OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=4/5, (7, -2) y=?

OpenStudy (anonymous):

y = mx+b y = 4/5(x) + b substitute (7,-2) in the equation and find b then y = mx+b is the line you are looking for

OpenStudy (amistre64):

y = (4/5)x -(4/5)(7) -2

OpenStudy (amistre64):

\[y = \frac{4x -38}{5}\]

OpenStudy (amistre64):

\[y=\frac{4}{5}x-\frac{38}{5}\]